OpenStudy (anonymous):

Which of the following represents fifth root of x squared in exponential form? x to the 2 fifths power x to the 5 over 2 power 5x2 2x5

OpenStudy (dumbcow):

\[\sqrt[n]{x^m} = x^{m/n}\] just remember the root is always the denominator of the fractional exponent

OpenStudy (anonymous):

that wasnt an answer though

OpenStudy (anonymous):

@driftracer305

OpenStudy (driftracer305):

yea so, just like \[(x)^m/n\]

OpenStudy (anonymous):

ok

OpenStudy (driftracer305):

it equals \[\sqrt[n]{(x)^m}\]

OpenStudy (driftracer305):

so \[\sqrt[5]{x^2}\]

OpenStudy (driftracer305):

it would be wat?

OpenStudy (anonymous):

im not sure

OpenStudy (driftracer305):

k it would be \[x^2/5\]

OpenStudy (driftracer305):

so A is correct.........

OpenStudy (anonymous):

ok thanks

OpenStudy (driftracer305):

@thedukelife .... glad to help buddy.......

OpenStudy (dumbcow):

critical thinking ... taking generalized concepts and applying them to specific problems try to associate similar forms, what do "m" and "n" represent ?

OpenStudy (driftracer305):

lol @dumbcow ....... dont take it too seriously... he just has some problems...... but u r right... he should've got it............ but were are here to help regardless....